# ColBERT Retrieval and Ensemble Response Scoring for Language Model Question Answering

### v1

@@ I Introduction
Advances in Large Language Models (LLMs) have significantly enhanced their performance across various Natural Language Processing (NLP) tasks. LLM series such as OpenAI’s GPTs [1, 2], have grown substantially in both model parameters and training data volume, driving much of this progress. One notable area of improvement is open-domain question answering (QA) [3]. Pre-trained on a wide range of domain-general data, LLMs excel in leveraging common knowledge [4]. However, this strength does not fully extend to domain-specific QA, which demands specialized knowledge that is often underrepresented in training datasets [4, 5]. Furthermore, LLMs can sometimes produce responses that conflict with established facts [6], posing a significant challenge in domain-specific QA where factual accuracy is crucial.
@@ II Related Work
The application of LLMs in highly specialized domains, such as telecommunications, presents unique challenges due to complex and dynamic documentation [13]. Ahmed et al. [10] evaluated the performance of LLMs in the telecommunications domain, focusing on zero-shot evaluations of models like Llama-2, Falcon, Mistral, and Zephyr. These models, more resource-efficient than ChatGPT, demonstrated comparable performance to fine-tuned models, underscoring the benefits of extensive pre-training. TelcoRAG [13], an open-source RAG framework, was developed to address the challenges in processing telecommunications standards, specifically from the 3GPP. This framework provides a foundation for applying RAG in telecommunications and other technical fields, offering valuable guidelines for deploying RAG pipelines on specialized content. The study [13] also provides a foundation for using LLMs for QA tasks for technical domain, particularly telecommunications.
@@ III-A Phi-2
For the Phi-2 track, we implemented a comprehensive system that integrates fine-tuning, retrieval-augmented generation, and domain-specific lexicon enhancement. Figure 1 illustrates the architecture of our Phi-2 QA system, demonstrating the interplay between these components.
@@ III-A2 Retrieval-Augmented Generation (RAG)
For the retrieval component, we used the ColBERT model [11]. ColBERT’s ability to capture nuanced semantic relationships is particularly beneficial in the highly technical telecommunication domain. The resulting vectorized chunks form the foundation of our retrieval system.
@@ III-A4 Prompt Engineering
  • Clear instructions for the task
@@ III-B Falcon-7B
For the Falcon-7B track, we adopt the same retrieval pipeline as used for Phi-2, the only difference being the number of chunks used as context. Unlike Phi-2 where we used as many chunks as the context window allowed, for Falcon-7B we found that providing more than 3 chunks in the prompt degrades the performance as shown in  Table I. We thus use 3 chunks in the final system. Following our Phi-2 approach, we also address the lexical mismatch by expanding the abbreviations found in the query and options. The key differences between the two systems are the prompts used and the scoring mechanism developed for the responses generated by Falcon-7B. The final system architecture is shown in Figure 1.
@@ III-B1 Prompting without Options
From our initial experiments on Falcon-7B we found that the model struggles to reason when presented with options for multiple choice QA. We found no apparent ability of the language model to use the question, context, and options to reason towards an informed response. Therefore, we do not provide the options in the prompt. With this change alone, we found that the LLM can use the context to generate a relevant answer to the query.
@@ III-B2 Response Scoring
To quickly test this hypothesis, we chose to use the cosine similarity between the embeddings of the response and the options as a measure of semantic similarity. We developed a fine-tuning objective for the embedding model that seeks to maximize the similarity between texts that entail one another, and minimize the similarity between contradicting texts. We select the large General Text Embedding (GTE) v1.5 model [18] and fine-tune it with triplet loss. Given the limited training set permitted for the challenge, we found that the best proxy for the responses seen at inference time to be the explanations for the correct options in the training set. For each question, we selected the explanation as the anchor, the correct option as the positive instance and a randomly selected option as the negative instance.
@@
The final score is a weighted sum of the term overlap and the cosine similarity. We run ablations for different static weights and select $\alpha_{1}=0.2$ and $\alpha_{2}=0.8$ for the scoring equation (1).
$score=\alpha_{1}\times overlap+\alpha_{2}\times similarity$ (1)
Where:
@@ IV-A Phi-2 Track: Retrieval Method Exploration and Parameter Sensitivity
Our hypothesis was that the ensemble approach could take advantage of the strengths of both sparse and dense retrieval, potentially yielding superior results in the highly technical telecommunication domain, particularly when working with the smaller Phi-2 model.
@@ IV-A3 BM25 and ColBERT Ensemble
Contrary to our initial hypothesis, the ensemble approach did not outperform the use of ColBERT alone with Phi-2. This finding provides valuable insights into the nature of our dataset and the effectiveness of dense retrieval methods in our specific context:
  • Lexical Similarity: The superior performance of ColBERT alone suggests that there wasn’t a significant lexical gap between the queries (questions) and the relevant documents in our dataset. A lexical gap would be expected if the queries used substantially different vocabulary or phrasing compared to the document content. The absence of this gap is likely due to our evaluation set being drawn from the same distribution as the training data, ensuring consistent terminology and phrasing.
  • Semantic Understanding: ColBERT’s effectiveness in our domain-specific task underscores its ability to capture nuanced semantic relationships, even when working with a smaller model like Phi-2. This suggests that semantic understanding was more crucial than handling lexical variations in our dataset.
@@
  • Context Quality vs. Quantity: As we increased the number of chunks retrieved from BM25 in our ensemble, we observed a deterioration in overall system performance. This suggests that the quality of retrieved context, particularly the semantic relevance captured by ColBERT, was more important than the quantity of keyword-matched chunks from BM25.
These findings reinforce the effectiveness of dense retrieval methods, particularly ColBERT, in capturing the nuanced semantic relationships present in our domain-specific corpus. However, it is important to note that while this approach was optimal for our current dataset, the dynamics might change in a real-world deployment scenario where end-users formulate their own queries. In such cases, a wider variety of lexical variations might be encountered, potentially making an ensemble approach more beneficial.
@@ IV-B1 Context Utilization
We evaluated the ability of Falcon-7B to utilize the entire context window, as was demonstrated by Phi-2. We selected the optimal chunk size determined from the Phi-2 experiment recorded in Figure 2 and increased the number of chunks added in the prompt. In  Table I it is observed that Falcon-7B is unable to disregard the noise introduced by irrelevant passages, or refer to relevant passages that are introduced later in the prompt.
TABLE I: Accuracy (%) for different top-k values of context provided in the Falcon-7B prompt
@@ IV-B2 Optimal Scoring Weights
To find the optimal weights for the term overlap and similarity scores, we uniformly varied the weights and scored the same set of responses. From the results in Table II we found that, in general, giving a higher weight to the similarity score is more beneficial. However, incorporating the overlap score with a small weight yields better accuracy than relying solely on the similarity score. Using the overlap score alone is the least effective, as it fails to capture the semantics needed to match the correct response to the appropriate option, which cannot be fully conveyed through shared terms alone.
TABLE II: Accuracy (%) for different scoring weights
@@ IV-B3 Enhancing the Embedding Model with an Entailment Objective
We evaluate the performance of the large GTE [18] embedding model prior to and following fine-tuning with the entailment objective. We report the accuracy values from scoring the responses with only the cosine similarity scores from the embeddings. The accuracy is $47.3\%$ when scoring with the base model, and $58.2\%$ with the fine-tuned model.
@@ IV-C Comparisons with leading LLMs
We assessed the performance of our QA systems in comparison to several top LLMs. In these experiments, the models were not fine-tuned. Table III (a) shows results for experiments where context, technical abbreviation expansions (TAEs), and options were included in the prompt. The highest accuracy was achieved by Llama 3.1 405B [21], outperforming our Phi-2 QA system by 6.3%. In Table III (b), we excluded context and TAEs while retaining the options to evaluate the models’ inherent knowledge. Here, we observe that the larger LLMs significantly outperform the smaller LLM. We hypothesize that this is largely attributed to the larger training datasets used when scaling LLMs and the additional capacity of the models to encode relevant information in their parameters. Finally, Table III (c) assesses the performance with context and TAEs included but without options. Llama 3.1 405B outperform the rest. The model appeared to be better than other LLMs at distilling their responses into short sentences that more closely resemble the correct options.
TABLE III: Accuracy (%) for different LLMs on the public subset of the challenge evaluation dataset
@@ IV-D Out of domain evaluation
To assess the generalizability of our methods in different domains, we evaluated our QA systems on the pharmacology subset of the MedMCQA [24] validation dataset. We selected pharmacology because it was one of the larger subsets for which we had access to a knowledge base of relevant textbooks. As presented in Table IV, the Phi-2 model fine-tuned on pharmacology questions achieved the highest accuracy. Notably, the Phi-2 model fine-tuned on telecommunication questions followed closely, with just a 1.7% difference. These findings imply that our fine-tuning objectives primarily enhance instruction and task alignment, rather than knowledge alignment. The Falcon-7B QA system did not outperform the baseline, underscoring the importance of constraining reasoning through the provided options. Interestingly, the optimal scoring weights for this evaluation were the same as those for the telecommunication questions, despite most options in MedMCQA being shorter than those in the challenge dataset. This further highlights the robustness of the ensemble scoring system in multiple-choice QA.
TABLE IV: Accuracy (%) on the pharmacology subset of the MedMCQA validation set
@@
V Conclusion
@@
VI Future Work
There are several promising avenues for future research that could further extend and refine our findings.
  • Explore dynamic context selection by resizing context based on query relevance, using a learnable scoring mechanism, potentially optimized via RLHF.
  • Improve coherence in retrieved chunks by exploring context stitching to maintain logical order.
  • Incorporate multi-modal input handling by adding document parsers that can process diagrams.
  • Implement sub-domain specific adapters for the language model, inspired by the MAD-X framework [26], to enhance handling of diverse telecommunications topics through dynamic adapter selection based on query classification.

### v2

@@ I Introduction
Advances in Large Language Models (LLMs) have significantly enhanced their performance across various Natural Language Processing (NLP) tasks. LLM series such as OpenAI’s GPTs [1, 2], have grown substantially in both model parameters and training data volume, driving much of this progress. One notable area of improvement is open-domain question answering (QA) [3]. Pre-trained on a wide range of domain-general data, LLMs excel in leveraging common knowledge [4]. However, this strength does not fully extend to domain-specific QA, which demands specialized knowledge often underrepresented in training datasets [4, 5]. Furthermore, LLMs can sometimes produce responses that conflict with established facts [6], posing a significant challenge in domain-specific QA where factual accuracy is crucial.
@@ II Related Work
The application of LLMs in highly specialized domains, such as telecommunications, presents unique challenges due to complex and dynamic documentation [13]. Ahmed et al. [10] evaluated the performance of LLMs in the telecommunications domain, focusing on zero-shot evaluations of models like Llama-2, Falcon, Mistral, and Zephyr. These models, more resource-efficient than ChatGPT, demonstrated comparable performance to fine-tuned models, underscoring the benefits of extensive pre-training. TelcoRAG [13], an open-source RAG framework, was developed to address the challenges in processing telecommunications standards, specifically from the 3GPP. This framework provides a foundation for applying RAG in telecommunications and other technical fields, offering valuable guidelines for deploying RAG pipelines on specialized content. The study [13] also provides a foundation for using LLMs for QA tasks in technical domain, particularly telecommunications.
@@ III-A Phi-2
For the Phi-2 track, we implemented a comprehensive system that integrates fine-tuning, RAG, and domain-specific lexicon enhancement. Figure 1 illustrates the architecture of our Phi-2 QA system, demonstrating the interplay between these components.
@@ III-A2 Retrieval-Augmented Generation (RAG)
For the retrieval component, we used the ColBERT model [11]. ColBERT’s ability to capture intricate meaning between terms is particularly beneficial in the highly technical telecommunication domain. The resulting vectorized chunks form the foundation of our retrieval system.
@@ III-A4 Prompt Engineering
  • Unambiguous instructions for the task
@@ III-B Falcon-7B
For the Falcon-7B track, we adopt the same retrieval pipeline as used for Phi-2, the only difference being the number of chunks used as context. Unlike Phi-2 where we used as many chunks as the context window allowed, for Falcon-7B we found that providing more than 3 chunks in the prompt degrades the performance as shown in  Table II. We thus use 3 chunks in the final system. Following our Phi-2 approach, we also address the lexical mismatch by expanding the abbreviations found in the query and options. The key differences between the two systems are the prompts used and the scoring mechanism developed for the responses generated by Falcon-7B. The final system architecture is shown in Figure 1.
@@ III-B1 Prompting without Options
From our initial experiments on Falcon-7B we found that the model struggles to reason when presented with options for multiple choice QA. We found no apparent ability of the language model to use the question, context, and options to reason toward an informed response. Therefore, we do not provide the options in the prompt. With this change alone, we found that the LLM can use the context to generate a relevant answer to the query.
@@ III-B2 Response Scoring
To quickly test this hypothesis, we chose to use the cosine similarity between the embeddings of the response and the options as a measure of semantic similarity. We developed a fine-tuning objective for the embedding model that seeks to maximize the similarity between texts that entail one another and minimize the similarity between contradicting texts. We select the large General Text Embedding (GTE) v1.5 model [18] and fine-tune it with triplet loss. Given the limited training set permitted for the challenge, we found that the best proxy for the responses seen at inference time to be the explanations for the correct options in the training set. For each question, we selected the explanation as the anchor, the correct option as the positive instance, and a randomly selected option as the negative instance.
@@
The final score is a weighted sum of the term overlap and the cosine similarity. We run ablations for different static weights and select $\alpha_{1}=0.2$ and $\alpha_{2}=0.8$, where:
@@ IV-A Phi-2 Track: Retrieval Method Exploration and Parameter Sensitivity
We hypothesized that the ensemble approach could take advantage of the strengths of both sparse and dense retrieval, potentially yielding superior results in the highly technical telecommunication domain, particularly when working with the smaller Phi-2 model.
@@ IV-A3 BM25 and ColBERT Ensemble
Contrary to our initial hypothesis, the ensemble approach did not outperform the use of ColBERT alone with Phi-2. To understand why, we investigated the relevance of the retrieved chunks for both methods. Given the labor-intensive nature of human evaluation, we initially employed an LLM to assess recall at 13 (the optimal number of chunks for Phi-2), prompting it to determine whether each passage was sufficient to answer a question. Using GPT-4o-mini [20], we found that for 96.99% of the questions in the public evaluation set, the correct answer appeared within the top 13 retrieved chunks. However, upon closer inspection of the results, we found this number to be grossly overestimated due to a significant number of false positives.
To obtain a more accurate estimate, we randomly sampled 122 questions from the public evaluation set and manually evaluated the binary recall for the ColBERT and BM25 retrieval systems. In  Table I we report the results of this evaluation.
TABLE I: Proportion of Questions where an Answer-Bearing Chunk was Found in the Top 13 Retrieved Chunks
Method Number of Questions Percentage of Questions
ColBERT 98 80.3%
BM25 94 77.1%
This finding provides valuable insights into the nature of our dataset and the effectiveness of dense retrieval methods in our specific context:
  • Lexical Similarity: The superior performance of ColBERT alone suggests that there was no significant lexical gap between the queries (questions) and the relevant documents in our dataset. A lexical gap would be expected if the queries used substantially different vocabulary or phrasing compared to the document content. The absence of this gap is likely due to our evaluation set being drawn from the same distribution as the training data, ensuring consistent terminology and phrasing.
  • Semantic Understanding: ColBERT’s effectiveness in our domain-specific task underscores its ability to model semantic relationships; it can recognize subtle distinctions between words and phrases, even when working with a smaller model like Phi-2. This suggests that semantic understanding was more crucial than handling lexical variations in our dataset.
@@
  • Context Quality vs. Quantity: As we increased the number of chunks retrieved from BM25 in our ensemble, we observed a deterioration in overall system performance. This indicates that the quality of retrieved context, particularly the semantic relevance captured by ColBERT, was more important than the quantity of keyword-matched chunks from BM25.
These findings reinforce the effectiveness of dense retrieval methods, particularly ColBERT, in capturing the intricate semantic patterns present in our domain-specific corpus. However, it is important to note that while this approach was optimal for our current dataset, the dynamics might change in a real-world deployment scenario where end-users formulate their own queries. In such cases, a wider variety of lexical variations may be encountered, potentially making an ensemble approach more beneficial.
@@ IV-B1 Context Utilization
We evaluated the ability of Falcon-7B to utilize the entire context window, as was demonstrated by Phi-2. We selected the optimal chunk size determined from the Phi-2 experiment (shown in Figure 2) and increased the number of chunks added to the prompt. In  Table II it is observed that Falcon-7B is unable to disregard the noise introduced by irrelevant passages, or refer to relevant passages that are introduced later in the prompt.
TABLE II: Accuracy (%) for different top-k values of context provided in the Falcon-7B prompt
@@ IV-B2 Optimal Scoring Weights
To determine the optimal weights for the term overlap and similarity scores, we uniformly varied the weights and scored the same set of responses. From the results in Table III, we found that giving a higher weight to the similarity score is generally more beneficial. However, incorporating the overlap score with a small weight yields better accuracy than relying solely on the similarity score. Using the overlap score alone is the least effective, as it fails to capture the semantics needed to match the correct response to the appropriate option, which cannot be fully conveyed through shared terms alone.
TABLE III: Accuracy (%) for different scoring weights
@@ IV-B3 Enhancing the Embedding Model with an Entailment Objective
We evaluate the performance of the large GTE [18] embedding model before and after fine-tuning it with the entailment objective. We report the accuracy values from scoring the responses with only the cosine similarity scores from the embeddings. The accuracy is $47.3\%$ when scoring with the base model, and $58.2\%$ with the fine-tuned model.
@@ IV-C Comparisons with leading LLMs
We assessed the performance of our QA systems in comparison to several top LLMs. In these experiments, the models were not fine-tuned. Table IV (a) shows results for experiments where context, technical abbreviation expansions (TAEs), and options were included in the prompt. The highest accuracy was achieved by Llama 3.1 405B [22], outperforming our Phi-2 QA system by 6.3%. In Table IV (b), we excluded context and TAEs while retaining the options to evaluate the models’ inherent knowledge. Here, we observe that the larger LLMs significantly outperform the smaller LLM. We hypothesize that this is largely attributed to the larger training datasets used when scaling LLMs and the additional capacity of the models to encode relevant information in their parameters. Finally, Table IV (c) assesses the performance with context and TAEs included but without options. Llama 3.1 405B outperform the rest. The model appeared to be better than other LLMs at distilling their responses into short sentences that more closely resemble the correct options.
TABLE IV: Accuracy (%) for different LLMs on the public subset of the challenge evaluation dataset
@@ IV-D Out of domain evaluation
To assess the generalizability of our methods across different domains, we evaluated our QA systems on the pharmacology subset of the MedMCQA [24] validation dataset. We selected pharmacology because it was one of the larger subsets for which we had access to a knowledge base of relevant textbooks. As presented in Table V, the Phi-2 model fine-tuned on pharmacology questions achieved the highest accuracy. Notably, the Phi-2 model fine-tuned on telecommunication questions closely followed, with just a 1.7% difference. These findings imply that our fine-tuning objectives primarily enhance instruction and task alignment, rather than knowledge alignment. The Falcon-7B QA system did not outperform the baseline, underscoring the importance of constraining reasoning through the provided options. Interestingly, the optimal scoring weights for this evaluation were the same as those for the telecommunication questions, despite most options in MedMCQA being shorter than those in the challenge dataset. This further highlights the robustness of the ensemble scoring system in multiple-choice QA.
TABLE V: Accuracy (%) on the pharmacology subset of the MedMCQA validation set
@@
V Analysis
V-A Inference Time
To complement our accuracy results, we conducted timing analyses for the complete pipeline execution on an NVIDIA L4 GPU. Table VI presents the execution times for both Phi-2 and Falcon-7B systems under various configurations. These measurements were taken at the optimal $k$ values for each system ($k=13$ for Phi-2, $k=3$ for Falcon-7B), representing the best trade-off between accuracy and inference time in our experiments.
TABLE VI: QA Pipeline Inference Time Results
Model Context Chunks ($k$) Inference Time (s)
Phi-2 13 2.102
Phi-2 0 0.524
Falcon-7B 3 5.023
Falcon-7B 0 4.682
The pipeline execution times reveal significant differences between the two systems. The Phi-2-based system demonstrates faster execution times despite using FP32 precision, compared to Falcon-7B’s FP16. This disparity is attributed to several factors:
  • Model size: Falcon-7B, being substantially larger, requires more computational resources for inference.
  • Token generation limits:
    • 10 tokens for Phi-2 - since we fine-tuned it to generate the option within the first few tokens.
    • 100 tokens for Falcon-7B - since the free-form response was often more verbose.
  • Context utilization: The addition of context chunks ($k$) increases execution time for both systems, with Phi-2 showing a more pronounced relative impact.
For the Falcon-7B system, we additionally measured the time for the ensemble scoring process, which adds a relatively small overhead of approximately 49 milliseconds. This step is unique to the Falcon-7B pipeline and contributes to its overall execution time.
Our approach introduces additional parameters beyond the core models. The embedding model for scoring in the Falcon-7B system adds 434.1M parameters, while the LoRA adapter for Phi-2 introduces 26.2M parameters. However, it’s worth noting that the LoRA adapter’s parameter count is relatively small compared to the base model size, and does not significantly impact the inference time. These additional parameters contribute to the overall system complexity but play crucial roles in enhancing performance and enabling domain-specific adaptations.
V-B Error Analysis
To categorize the errors made by our Phi-2 QA system, we calculated the error rates for each dataset category (standards specifications and standards overview). Here, the error set refers to the set of questions from the public evaluation set that our Phi-2 system answered incorrectly. We expected that a well-represented category with a notable relative increase in its error set distribution, compared to its dataset distribution, would indicate areas where the model struggles. However, as seen in Table VII, the observed differences are not significant enough to suggest that any particular category presents a consistent challenge for the model.
TABLE VII: Frequencies per Category: Dataset vs Error Set
Standards Overview Standards Specifications
Dataset Count 63 303
Error Set Count 9 65
Dataset % 17.2% 82.8%
Error Set % 12.2% 87.8%
Relative % Change -29.3% 6.1%
Stronger conclusions can be drawn from the number of questions for which at least one relevant answer-bearing passage was retrieved. As shown in Table VIII, for half of the questions where the Phi-2 QA system failed, no relevant passage was found. This suggests that the main bottleneck in our system is the retrieval pipeline.
TABLE VIII: Frequencies for Binary Recall: Dataset vs Error Set
0 1
Dataset Count1 24 98
Error Set Count 37 37
Dataset % 19.7% 80.3%
Error Set % 50.0% 50.0%
Relative Change (%) 154.2% -37.8%
  • 1
    Here, dataset refers to the randomly selected subset of questions for which binary recall was human evaluated.
VI Conclusion
@@
VII Future Work
There are several promising avenues for future research that could further extend and refine our findings:
  • Conduct a more comprehensive analysis of the retrieval pipeline:
    • Extend our preliminary binary precision analysis at $k=13$ to various $k$ values.
    • Evaluate additional metrics such as mean reciprocal rank, precision, and recall at different $k$ values to gain deeper insights into retrieval performance.
  • Explore dynamic context selection by resizing context based on query relevance, using a learnable scoring mechanism, potentially optimized via Reinforcement Learning from Human Feedback (RLHF).
  • Improve coherence in retrieved chunks by exploring context stitching techniques to maintain logical order and enhance the quality of the provided context.
  • Incorporate multi-modal input handling by developing document parsers capable of processing diagrams, thereby expanding the range of information our system can utilize.